On December 31, 2010, Merlin Company had outstanding 400,000 shares of common stock and 40,000 shares of 8% cumulative preferred stock (par $10). On February 1, 2011, Merlin issued an additional 35,000 shares of common stock. A 15% stock dividend was declared and distributed on July 1, 2011. On September 30, 2011, 9,000 shares were retired. At year-end, there were fully vested incentive stock options outstanding for 30,000 shares of common stock (adjusted for the stock dividend). The exercise price was $18. The market price of the common stock averaged $20 during the year. Also outstanding were $1,000,000 face amount of 10% convertible bonds issued in 2008 and convertible into 50,000 common shares (adjusted for the stock dividend). Net income was $900,000. The tax rate for the year was 40%.
Compute basic and diluted EPS for the year ended December 31, 2011.

show work
Answer #1
Basic EPS = Net income-preferred dividend /weighted avg commonoutstanding share
Preferred Dividend = 400000*8% = 32000
Weighted average common outstanding share
(400000*1.15)+(35000*11/12)*1.15-(9000*3/12)
494646
Weighted avg common outstanding share = 494646
Basic EPS = (900000-32000)/494646
Basic EPS = 1.75
Calculation of Diluted EPS
adjusted net income
900000-32000+(1000000*10%)-(100000*40%)
adjusted net income = 928000
Ajusted weighted average common outstanding share
400000*1.15+(35000*11/12)*1.15- (9000*3/12)+(30000-27000**)+50000
adjusted weighted average common outstanding share= 547646
Diluted EPS = 928000/547646 = 1.69
** 30000 shares*18 (exercise price) =540000
540000/20 avg market price =27000